# Quantifying the Impact of the Si/O Interface in CCSN Explosions Using the Force Explosion Condition

### v1

@@ Abstract
pubyear: 2024pagerange: Quantifying the Impact of the Si/O Interface in CCSN Explosions Using the Force Explosion ConditionQuantifying the Impact of the Si/O Interface in CCSN Explosions Using the Force Explosion Condition
@@ 2.1 The generalized Force Explosion Condition
$\tilde{L}_{\nu}\tau_{g}+\tilde{W}_{b}-a\tilde{\kappa}+c\left\langle\tilde{R}^{r}_{r}\right\rangle=b.$ (2)
@@ 2.3 Numerical setup
The spatial resolution was the same for all of the simulated progenitors (with the only exception of $<10$ progenitors of very low compactness, with $M_{\rm ZAMS}<11$). We adopted a grid of 700 zones linearly spaced up to 20 km, with a resolution of 300 m, and then logarithmically spaced out to 15,000 km. The decision of using the same spatial grid for all progenitors was motivated by the fact that the extra heating from STIR has a weak dependence on the spatial resolution, as shown in Boccioli et al. (2022), particularly behind the shock.
@@ 3 Explosion dynamics of 341 progenitors
One of the primary aims of this manuscript is to use FEC+ to understand why the Si/O interface facilitates an explosion for most 1D+ STIR simulations and why the Si/O interface does not trigger explosions for other progenitors. The simulations for this study are described in Boccioli & Fragione (2024). In particular, Figure 1 shows the evolution of the shock radius and the FEC+ for 341 GR1D+ simulations of KEPLER progenitors from Sukhbold et al. (2016) and Woosley et al. (2002), with masses ranging from 9 $M_{\odot}$ to 120 M and metallicities of $z=0$, $z=10^{-4}z_{\odot}$, and $z=z_{\odot}$, where $z_{\odot}$ indicates solar metallicity. The solid-green curves show simulations that eventually explode, and the black curves shows simulations that fail to explode. The most striking feature in the FEC+ plot (right panel) is that there is a clear separation between failed and successful explosions for a value of FEC+ around 0.28–0.3. This is yet another example showing that the FEC+ can be used as an explosion diagnostic. In this manuscript, we will use the FEC+ to diagnose why the Si/O interface helps facilitate explosions sometimes but not always.
@@ 3.1 Pre-explosion phase in light of the FEC+
The remaining $\sim 3-4\%$ of the analyzed simulations are outliers. For example, four progenitors with large compactness do not explode, whereas 10 progenitors with large Si/O interfaces accreted towards the beginning or the end of the stalled-shock phase do not explode. It is worth noting here that while these outliers do not conform to the Si/O interface condition for explosion, all simulations are consistent with the FEC+ explosion condition, i.e. the FEC+ never crosses the threshold. Perhaps the most interesting outlier is the 24.2 $M_{\odot}$ progenitor with $z=10^{-4}z_{\odot}$, that accretes a large density drop very late (and therefore is expected to fail) but instead leads to a successful explosion. We include this progenitor in Figure 2 to highlight the unusual time evolution of the FEC+ and shock radius. Two more progenitors, i.e. the 23.4 and 23.2 $M_{\odot}$ with $z=10^{-4}z_{\odot}$ have a similar pre-SN structure and post-bounce evolution. However, these progenitors have high compactness ($\xi_{2.0}>0.5$), and the Si/O interface is accreted a bit earlier. Therefore their shock radius and FEC+ time evolution are not quite as unusual as they are for the 24.2 $M_{\odot}$ progenitor. In other words, the FEC+ is not far from the threshold when the explosion sets in, whereas it clearly is for the 24.2 $M_{\odot}$ progenitor.
@@
The 19.5 M progenitor shows a unique behavior. A very large density drop is accreted early in the evolution. However, since the FEC+ is still small when the accretion occurs, it does not increase enough to go above the threshold. Therefore, the explosion is not triggered after the accretion of the interface, and in principle, one would expect the explosion to fail. However, since this early accretion causes the shock to stall at larger radii than any of the other progenitors, the FEC+ at around $150-200$ ms is the largest among the progenitors shown. Therefore, between 200 and 300 ms, the FEC+ is barely below the threshold. Eventually, at around 600-700 ms, the FEC+ finally goes above the threshold causing a (relatively) late explosion. One can speculate about what the cause of the explosion is. For most progenitors, it is the accretion of the Si/O interface that directly launches the explosion by reducing the ram pressure on the shock. For this specific progenitor (as well as a handful of others), the explosion happens much later than the accretion.
@@
Nonetheless, the early accretion of the Si/O interface still plays a crucial role in causing the explosion. This can be seen by comparing the shock radius and FEC+ evolution for the 19.5 M and the 15.5 M progenitors. These have almost identical pre-SN density (and also compositional) structures, as seen in the third panel of Figure 2, with the important difference that the 15.5 M progenitor has a smaller density drop at the Si/O interface, and also a Si shell that is 0.125 M thicker, causing the interface to be accreted at a slightly later time. The smaller density drop causes the shock radius and the FEC+ to have a smaller increase, and therefore the FEC+ during the stalled-shock phase (i.e. at $\sim 200$ ms after bounce) is smaller for the 15.5 M progenitor. This is the reason why the 15.5 M progenitor fails whereas the 19.5 M progenitor explodes at a much later time.
@@
Another interesting feature of these two progenitors is that, since they have a very similar post-bounce evolution, the shock trajectory of the 15.5 M progenitor would look pretty much exactly like the one for the 19.5 M progenitor if one were to use a larger value of $\alpha_{\rm MLT}$. This shows that most CCSNe are in general very close to explosion, and small changes in the physics can lead to qualitatively drastically different outcomes. In the next section, we will investigate how changes in the strength of $\nu$-driven convection (i.e. changes in $\alpha_{\rm MLT}$) and in the density drop at the Si/O interface can change the shock and the evolution of the FEC+, leading to a successful or failed explosion. It is important to remark that also changes in the nuclear equation of state, neutrino opacities, asymmetries in the pre-SN progenitor, beyond standard model physics, collective neutrino oscillations, and more in general any physical input of the simulation, can lead to similar differences in the outcome of the supernova.
@@
Figure 3: The top (bottom) panel shows the evolution of the shock radius (FEC+) for a 20 $M_{\odot}$ MESA progenitor (Farmer et al., 2016) whose collapse was simulated using GR1D+ for different values of $\alpha_{\rm MLT}$, shown in the colorbar on the right. Before calculating the value of the FEC+, each quantity in eq. (2) has been smoothed with a Savitzky-Golay filter in a 10 ms window. The filled dots show the time before the accretion of the Si/O interface, whereas the filled diamonds show the time after the accretion of the Si/O interface. All of the simulations with $\alpha_{\rm MLT}\geq 1.495$ successfully explode, and the accretion of the Si/O interface leads to a larger increase of the FEC+ for larger values of $\alpha_{\rm MLT}$. To avoid cluttering, we only show exploding simulations with $\alpha_{\rm MLT}$ spaced by 0.01, but a finer grid is shown in Figures 4 and 5.
@@ 4 Interplay between neutrino-driven convection and accretion of the Si/O interface
As shown in the previous section, the accretion of the Si/O interface can often lead to the explosion of a CCSN. In this section, we use the FEC+ to quantify in detail how the post-accretion expansion of the shock depends on both the strength of $\nu$-driven convection and the density drop at the Si/O interface. In addition to being a condition for explosion, the FEC+ also quantifies a distance from explosion. Therefore, it is useful as a quantitative measure of how various physics affect the explosion outcome. It should be highlighted that, in a more realistic 3D simulation, other multi-dimensional effects are also at play, and therefore the post-bounce phase is more complicated. However, the simplicity of 1D+ simulations enables systematic study of the interplay between the accretion of the Si/O interface and parametrized $\nu$-driven convection.
@@ 4.1 Sensitivity to the strength of neutrino-driven convection
For the accretion of the Si/O interface to be sufficient to launch the explosion, the FEC+ has to be already close to the threshold. This is shown in the top panel of Figure 4, where the x-axis is the value of the FEC+ at $t=t_{\rm accr}^{\rm start}$ (i.e. the circles in Figure 3) and the y-axis is the value of the FEC+ at $t=t_{\rm accr}^{\rm end}$ (i.e. the diamonds in Figure 3). Simulations with a value of FEC+ at $t=t_{\rm accr}^{\rm start}$ below $\sim 0.24$ do not explode, because once they accrete the interface the change in FEC+ is not enough to push it above threshold.
@@
A similar correlation can be seen between $\Delta{\rm FEC+}$ (or $\Delta R_{\rm shock}$) and the value of FEC+ (or shock radius) at $t_{\rm accr}^{\rm start}$ (and also $\alpha_{\rm MLT}$), as shown in Figure 5. The quantities $\Delta{\rm FEC+}$ and $\Delta R_{\rm shock}$ are defined as the differences between the values of FEC+ and shock radius after and before the accretion of the Si/O interface. Larger $\alpha_{\rm MLT}$, i.e. stronger $\nu$-driven convection, leads to the shock stalling at a larger radius, causing the FEC+ to be larger, and eventually the accretion of the Si/O interface causes a proportionally larger increase in FEC+ and shock radius.

### v2

@@ Abstract
pubyear: 2025pagerange: Quantifying the Impact of the Si/O Interface in CCSN Explosions Using the Force Explosion ConditionQuantifying the Impact of the Si/O Interface in CCSN Explosions Using the Force Explosion Condition
@@ 2.1 The generalized Force Explosion Condition
Notice that in this work we use the net neutrino heating rate in the gain region per unit time $\tilde{\dot{Q}}_{\nu}$ in place of the first term in Eq. (1) $\tilde{L}_{\nu}\tau_{g}$, as done in Gogilashvili et al. (2023b) when using self-consistent simulations to calculate the FEC. The reason behind this is that in the original semi-analytical FEC formulation, the term $L_{\nu}\tau_{g}$ is meant to represent the net neutrino heating rate in the gain region $\dot{Q}_{\nu}$, which can however be directly calculated from our simulations.
@@
$\tilde{\dot{Q}}_{\nu}+\tilde{W}_{b}-a\tilde{\kappa}+c\left\langle\tilde{R}^{r}_{r}\right\rangle=b.$ (2)
@@ 2.3 Numerical setup
The spatial resolution was the same for all of the simulated progenitors (with the only exception of $<10$ progenitors of very low compactness, with $M_{\rm ZAMS}<11$). We adopted a grid of 700 zones linearly spaced up to 20 km, with a resolution of 300 m, and then logarithmically spaced out to 15,000 km. The decision to use the same spatial grid for all progenitors was motivated by the fact that the extra heating from STIR has a weak dependence on the spatial resolution, as shown in Boccioli et al. (2022), particularly behind the shock.
@@ 3 Explosion dynamics of 341 progenitors
One of the primary aims of this manuscript is to use FEC+ to understand why the Si/O interface facilitates an explosion for most 1D+ STIR simulations and why the Si/O interface does not trigger explosions for other progenitors. The simulations for this study are described in Boccioli & Fragione (2024). In particular, Figure 1 shows the evolution of the shock radius and the FEC+ for 341 GR1D+ simulations of KEPLER progenitors from Sukhbold et al. (2016) and Woosley et al. (2002), with masses ranging from 9 $M_{\odot}$ to 120 M and metallicities of $z=0$, $z=10^{-4}z_{\odot}$, and $z=z_{\odot}$, where $z_{\odot}$ indicates solar metallicity. The solid-green curves show simulations that eventually explode, and the black curves show simulations that fail to explode. The most striking feature in the FEC+ plot (right panel) is that there is a clear separation between failed and successful explosions for a value of FEC+ around 0.28–0.3. This is yet another example showing that the FEC+ can be used as an explosion diagnostic.
Although the definition of a successful explosion can vary, in this manuscript, we define a successful shock revival when the shock crosses 500 km. We will be using the same definition for the explosion time in Section 5. All of the simulations that show a successful shock revival also have positive diagnostic energies $E^{+}_{\rm diag}$, where
[…]
where $e_{\rm bind}$ is the binding energy of the material and the integral is performed only over regions with positive binding energy (Buras et al., 2006; Bruenn et al., 2009). Notice that we subtracted the binding energy (i.e. thermal energy minus gravitational energy) of the yet unshocked material ($E_{\rm ov}$) to have a more realistic estimate of the final explosion energy. The explosion energies at the end of the simulations of the 341 progenitors described above are shown in Figure 3. Given the relatively small radial domain of the simulations, the shock leaves the boundary before the diagnostic energy has plateaued, especially in the case of high-compactness progenitors, and therefore Figure 3 represents a lower limit to the final explosion energy.
In 1D simulations, the mass accretion is shut off once the shock is revived, and therefore neutrino heating drops to zero, even though a central black hole never forms. However, multi-dimensional simulations show that continued accretion after shock revival can be quite significant. This causes neutrino heating to be nonzero even after shock revival, potentially leading to larger explosion energies compared to 1D simulations (Burrows et al., 2024). However, in the case of high-compactness progenitors, continued accretion causes an early collapse to a black hole, completely shutting off neutrino emission that, in some cases (Powell et al., 2021; Sykes & Müller, 2024; Eggenberger Andersen et al., 2024), can lead to a decrease of the final explosion energy. Overall, one can still see in Figure 3 that the explosion energy in 1D+ simulations increases with the compactness of the progenitor, as also seen in multi-dimensional simulations (Burrows et al., 2018).
It is also instructive to analyze the ratio of non-exploding models compared to observational constraints. To do that, we assumed that stars below 11 $M_{\odot}$ (i.e. the lowest mass simulated for the $z=0$ and $z=10^{-4}z_{\odot}$ progenitors) would produce a successful explosion. Then, we weighted our simulations based on their initial ZAMS mass using a Salpeter initial mass function with exponent $-2.35$ (Salpeter, 1955), independently of metallicity. We calculated the fraction of failed SNe to be $f_{\rm fSNe}=0.25-0.3$, depending on whether we assume the lowest mass of a star that explodes as a CCSN to be 8 $M_{\odot}$ or $9M_{\odot}$, respectively (see Figure 6 in Boccioli & Fragione (2024)). This is consistent with the observed fraction of failed supernovae $f_{\rm fSNe}=0.16^{+0.23}_{-0.12}$ (Neustadt et al., 2021).
In the remainder of this paper, we will use the FEC+ to diagnose why the Si/O interface can help facilitate explosions in many, but not all, successful explosions.
@@
Figure 3: Explosion energy calculated at the end of the simulations, as defined by Eq. (6). This represents a lower limit of the explosion energy, since when the shock leaves the computational domain the diagnostic energy is still increasing. As expected the explosion energy increases with the progenitor compactness. The apparent trends of lower metallicity yielding large explosion energies for high-compactness progenitors, however, might be due to the aforementioned small computational domain preventing the diagnostic energy to plateau. Therefore, simulations with a larger radial domain (currently underway) are needed before drawing robust conclusions regarding the difference in explosion energy for progenitors of different metallicities.
@@ 3.1 Pre-explosion phase in light of the FEC+
The remaining $\sim 3-4\%$ of the analyzed simulations are outliers. For example, four progenitors with large compactness do not explode, whereas 10 progenitors with large Si/O interfaces accreted towards the beginning or the end of the stalled-shock phase do not explode. It is worth noting here that while these outliers do not conform to the Si/O interface condition for explosion, all simulations are consistent with the FEC+ explosion condition, i.e. the FEC+ never crosses the threshold. Perhaps the most interesting outlier is the 24.2 $M_{\odot}$ progenitor with $z=10^{-4}z_{\odot}$, which accretes a large density drop very late (and therefore is expected to fail) but instead leads to a successful explosion. We include this progenitor in Figure 2 to highlight the unusual time evolution of the FEC+ and shock radius. Two more progenitors, i.e. the 23.4 and 23.2 $M_{\odot}$ with $z=10^{-4}z_{\odot}$ have a similar pre-SN structure and post-bounce evolution. However, these progenitors have high compactness ($\xi_{2.0}>0.5$), and the Si/O interface is accreted a bit earlier. Therefore their shock radius and FEC+ time evolution are not quite as unusual as they are for the 24.2 $M_{\odot}$ progenitor. In other words, the FEC+ is not far from the threshold when the explosion sets in, whereas it clearly is for the 24.2 $M_{\odot}$ progenitor.
@@
The 19.5 M progenitor shows a unique behavior. A very large density drop is accreted early in the evolution. However, since the FEC+ is still small when the accretion occurs, it does not increase enough to go above the threshold. Therefore, the explosion is not triggered after the accretion of the interface, and in principle, one would expect the explosion to fail. However, since this early accretion causes the shock to stall at larger radii than any of the other progenitors, the FEC+ at around $150-200$ ms is the largest among the progenitors shown. Therefore, between 200 and 300 ms, the FEC+ is barely below the threshold. Eventually, at around 600-700 ms, the FEC+ finally goes above the threshold causing a (relatively) late explosion.
@@
One can speculate about what the cause of the explosion is. For most progenitors, it is the accretion of the Si/O interface that directly launches the explosion by reducing the ram pressure on the shock. For this specific progenitor (as well as a handful of others), the explosion happens much later than the accretion. Nonetheless, the early accretion of the Si/O interface still plays a crucial role in causing the explosion. This can be seen by comparing the shock radius and FEC+ evolution for the 19.5 M and the 15.5 M progenitors. These have almost identical pre-SN density (and also compositional) structures, as seen in the third panel of Figure 2, with the important difference that the 15.5 M progenitor has a smaller density drop at the Si/O interface, and also a Si shell that is 0.125 M thicker, causing the interface to be accreted at a slightly later time. The smaller density drop causes the shock radius and the FEC+ to have a smaller increase, and therefore the FEC+ during the stalled-shock phase (i.e. at $\sim 200$ ms after bounce) is smaller for the 15.5 M progenitor. This is the reason why the 15.5 M progenitor fails whereas the 19.5 M progenitor explodes at a much later time.
@@
Another interesting feature of these two progenitors is that, since they have a very similar post-bounce evolution, the shock trajectory of the 15.5 M progenitor would look pretty much exactly like the one for the 19.5 M progenitor if one were to use a larger value of $\alpha_{\rm MLT}$. This shows that most CCSNe are in general very close to explosion, and small changes in the physics can lead to qualitatively drastically different outcomes.
In order to better illustrate the explosion dynamics of 1D+ models, we show in Figure 4 the average energies and luminosities for both a 1D and a 1D+ simulation of a 20 $M_{\odot}$ progenitor from Sukhbold et al. (2016), i.e. the same shown in Figure 2. The neutrino quantities for electron neutrinos and antineutrinos are quite similar in the 1D and 1D+ simulations since they are mostly determined by post-bounce mass accretion history, which is the same in 1D and 1D+ until the explosion sets in at around $\sim 220$ ms in the 1D+ model, causing mass accretion to stop and therefore neutrino energies and luminosities to decrease. In the case of heavy-lepton neutrinos, however, both average energies and luminosities are instead larger in the 1D+ model after the explosion sets in. This can be explained by the fact that heavy-lepton neutrinos are produced deeper in the PNS, and therefore are much less dependent on the accretion history of the progenitor (O’Connor, 2015).
The increase in the luminosity of heavy-lepton neutrinos at late times in 1D+ models can be attributed to PNS convection, which pushes the neutrinosphere at larger radii, increasing the area of emission, and therefore the luminosity. The reason for the larger neutrino energies in 1D+ simulations is instead more subtle. Convection in the PNS smooths out the density gradient, and therefore deep in the PNS, between $\sim 10^{13}$ and $10^{14}\ {\rm g/cm^{3}}$, the density gradient in the 1D case is shallower, whereas near the neutrinosphere, between $\sim 10^{12}$ and $10^{13}\ {\rm g/cm^{3}}$, it is steeper. The shallower gradient deep inside the PNS in the 1D simulation causes larger energy losses due to inelastic scattering, explaining the lower neutrino energies. This effect is subdominant for electron-flavor neutrinos since beta processes dominate the opacity.
A heuristic condition often used to determine whether an explosion sets in is based on the ratio of heating and advection timescales (Fernández, 2012; Buras et al., 2006; Marek & Janka, 2009; Murphy & Burrows, 2008; Janka, 2001; Thompson et al., 2005). A few different definitions of these two quantities are possible, and we adopt the following:
$\tau_{\rm adv}=\frac{M_{\rm gain}}{\dot{M}}~{},\qquad\tau_{\rm heat}=\frac{E_{\rm gain}}{\dot{\mathcal{Q}}_{\nu}}$ (8)
where $E_{\rm gain}$, $M_{\rm gain}$, and $\dot{M}$ are the thermal energy, total mass, and average mass accretion rate in the gain region. It has been proposed that successful shock revival occurs when $\tau_{\rm adv}/\tau_{\rm heat}>1$ (Thompson et al., 2005). Note that some groups Summa et al. (2018) define $E_{\rm gain}$ as the total energy in the gain region (i.e. thermal plus kinetic plus gravitational) instead of the thermal energy only. However, in our case using the total energy results in the ratio $\tau_{\rm adv}/\tau_{\rm heat}$ to increase above 1 slightly earlier (and in some cases much earlier) than when shock revival occurs. This ambiguity is part of the reason why a more fundamental condition like the FEC+ should be preferred. We show the ratio of these two timescales in Figure 5, where one can also see the difference in adopting STIR (solid lines) compared to a simple 1D simulation (dashed lines). Essentially, convection allows the shock to stall at a larger radius, causing neutrino heating to increase (i.e. $\tau_{\rm heat}$ to decrease) driving $\tau_{\rm adv}/\tau_{\rm heat}$ above 1. A more detailed study on how convection enables neutrino heating to increase is currently underway and is beyond the scope of this paper.
Figure 4: The top panel shows the total net neutrino heating in the gain region for the set of progenitors discussed in Section 3.1. The bottom panel shows the ratio of the advection and heating timescales. Solid and dashed lines indicate simulations with and without STIR, respectively. Circles mark the times when the Si/O interface is accreted through the shock for the simulations with STIR. To avoid clutter, and since it occurs essentially at the same time, we did not mark the time of accretion of the Si/O interface for the simulations without STIR. The ratio of timescales, as expected, has some similarities to the FEC+ shown in the middle panel of Figure 2. However, as discussed in the text the FEC+ is a more robust, less ambiguous condition.
Figure 5: The top panel shows the total net neutrino heating in the gain region for the set of progenitors discussed in Section 3.1. The bottom panel shows the ratio of the advection and heating timescales. Solid and dashed lines indicate simulations with and without STIR, respectively. Circles mark the times when the Si/O interface is accreted through the shock for the simulations with STIR. To avoid clutter, and since it occurs essentially at the same time, we did not mark the time of accretion of the Si/O interface for the simulations without STIR. The ratio of timescales, as expected, has some similarities to the FEC+ shown in the middle panel of Figure 2. However, as discussed in the text the FEC+ is a more robust, less ambiguous condition.
In the next section, we will investigate how changes in the strength of $\nu$-driven convection (i.e. changes in $\alpha_{\rm MLT}$) and in the density drop at the Si/O interface can change the shock and the evolution of the FEC+, leading to a successful or failed explosion. It is important to remark that also changes in the nuclear equation of state, neutrino opacities, asymmetries in the pre-SN progenitor, beyond standard model physics, collective neutrino oscillations, and more in general any physical input of the simulation, can lead to similar differences in the outcome of the supernova.
@@
Figure 6: The top (bottom) panel shows the evolution of the shock radius (FEC+) for a 20 $M_{\odot}$ MESA progenitor (Farmer et al., 2016) whose collapse was simulated using GR1D+ for different values of $\alpha_{\rm MLT}$, shown in the colorbar on the right. Before calculating the value of the FEC+, each quantity in eq. (2) has been smoothed with a Savitzky-Golay filter in a 10 ms window. The filled dots show the time before the accretion of the Si/O interface, whereas the filled diamonds show the time after the accretion of the Si/O interface. All of the simulations with $\alpha_{\rm MLT}\geq 1.495$ successfully explode, and the accretion of the Si/O interface leads to a larger increase of the FEC+ for larger values of $\alpha_{\rm MLT}$. To avoid cluttering, we only show exploding simulations with $\alpha_{\rm MLT}$ spaced by 0.01, but a finer grid is shown in Figures 7 and 8.
@@ 4 Interplay between neutrino-driven convection and accretion of the Si/O interface
As shown in the previous section, the accretion of the Si/O interface can often lead to the explosion of a CCSN. In this section, we use the FEC+ to quantify in detail how the post-accretion expansion of the shock depends on both the strength of $\nu$-driven convection and the density drop at the Si/O interface. In addition to being an explosion condition, the FEC+ can also quantify a distance from explosion. Therefore, it is useful as a quantitative measure of how various physics affect the explosion outcome. It should be highlighted that, in a more realistic 3D simulation, other multi-dimensional effects are also at play, and therefore the post-bounce phase is more complicated. However, the simplicity of 1D+ simulations enables a systematic study of the interplay between the accretion of the Si/O interface and parametrized $\nu$-driven convection.
@@ 4.1 Sensitivity to the strength of neutrino-driven convection
For the accretion of the Si/O interface to be sufficient to launch the explosion, the FEC+ has to be already close to the threshold. This is shown in the top panel of Figure 7, where the x-axis is the value of the FEC+ at $t=t_{\rm accr}^{\rm start}$ (i.e. the circles in Figure 6) and the y-axis is the value of the FEC+ at $t=t_{\rm accr}^{\rm end}$ (i.e. the diamonds in Figure 6). Simulations with a value of FEC+ at $t=t_{\rm accr}^{\rm start}$ below $\sim 0.24$ do not explode, because once they accrete the interface the change in FEC+ is not enough to push it above the threshold.
@@
A similar correlation can be seen between $\Delta{\rm FEC+}$ (or $\Delta R_{\rm shock}$) and the value of FEC+ (or shock radius) at $t_{\rm accr}^{\rm start}$ (and also $\alpha_{\rm MLT}$), as shown in Figure 8. The quantities $\Delta{\rm FEC+}$ and $\Delta R_{\rm shock}$ are defined as the differences between the values of FEC+ and shock radius after and before the accretion of the Si/O interface. Larger $\alpha_{\rm MLT}$, i.e. stronger $\nu$-driven convection, leads to the shock stalling at a larger radius, causing the FEC+ to be larger, and eventually, the accretion of the Si/O interface causes a proportionally larger increase in FEC+ and shock radius.